# WeFT: Weighted Entropy-driven Fine-Tuning for dLLMs

### v1

@@ Abstract
Diffusion models have recently shown strong potential in language modeling, offering faster generation compared to traditional autoregressive approaches. However, applying supervised fine-tuning (SFT) to diffusion models remains challenging, as they lack precise probability estimates at each denoising step. While the diffusion mechanism enables the model to reason over entire sequences, it also makes the generation process less predictable and often inconsistent. This highlights the importance of controlling key tokens that guide the direction of generation. To address this issue, we propose WeFT, a weighted SFT method for diffusion language models, where tokens are assigned different weights based on their entropy. Derived from diffusion theory, WeFT delivers substantial gains: training on s1K, s1K-1.1, and 3k samples from open-r1, it achieves relative improvements of 39%, 64%, and 83% over standard SFT on four widely used reasoning benchmarks (Sudoku, Countdown, GSM8K, and MATH-500). The code and models will be made publicly available.
@@ 1 Introduction
Recently, the landscape of large language modeling has been reshaped by the emergence of diffusion-based approaches  (Nie et al., 2025; Ou et al., 2025; Yang et al., 2025), which offer a fundamentally different perspective on sequence generation compared to autoregressive (AR) methods  Radford et al. (2018; 2019); Brown et al. (2020). Rather than relying on strictly sequential decoding, diffusion large language models (dLLMs) employ iterative refinement procedures that enable efficient parallel generation and open new possibilities for scaling language models.
@@
Specifically, when designing the SFT objective, (Nie et al., 2025) model the generation process of dLLMs as a continuous-time diffusion process, where the diffusion timestep $t$ ranges within $[0,1]$. Here, $t=0$ denotes a completely original, non-masked sequence, and $t=1$ denotes a fully masked sequence. As illustrated in Figure 1(a), let the original response of the model be $x_{0}$. At a given timestep $t$, each token is masked with probability $t$, resulting in a partially masked response $x_{t}$. The training objective of dLLMs is then to enable the model to reconstruct the masked parts of $x_{0}$ given the partially masked input $x_{t}$. Based on this formulation, the loss is defined as follows:
$L=\sum_{i}\mathbb{E}_{t}\left[\mathbf{1}[x_{t}^{i}=\mathbf{M}]\,\frac{1}{t}\log\big(x_{0}^{i}\mid x_{t}\big)\right]$ (1)
@@
To address this discrepancy, we introduce WeFT, a token-aware weighting mechanism based on predictive uncertainty. Specifically, we find that the entropy of the model’s predictive distribution over tokens serves as a reliable metric for token importance at the current training stage. Tokens associated with higher entropy are typically those that relate to reasoning or planning, and they also correspond to positions where the model exhibits greater uncertainty in generation. By prioritizing these high-entropy tokens during training via an adaptive weighting strategy applied to the SFT loss, we can encourage the model to allocate more capacity to the parts of the sequence that are crucial for planning and reasoning.
Figure 1: (a) The SFT pipeline. A timestep $t$ is uniformly sampled from $[0,1]$, and each token is masked independently with probability $t$. The training objective is to predict the masked tokens accurately based on the unmasked ones. (b) The WeFT pipeline. In each training step, we perform two forward passes. During the first forward pass, we mask the entire answer and estimate the masking rate $\beta_{i}$ for each token by computing its predictive entropy. In the second forward pass, the $i$-th token is masked with probability $t_{i}$ (computed from $\beta_{i}$), and its training weight is set to $\tfrac{1}{t_{i}}$. Tokens with higher entropy are more likely to be masked and thus receive stronger training signals.
Based on the fundamental principles of diffusion models, we derive a weighted SFT loss formulation that is theoretically consistent with diffusion dynamics. Empirically, we evaluate this approach on four widely used benchmarks, Sudoku (Arel, 2025), Countdown (Pan et al., 2025), GSM8K (Cobbe et al., 2021), and MATH500 (Lightman et al., 2024), by comparing models fine-tuned with the standard SFT algorithm against those fine-tuned with WeFT under identical training data. Specifically, our method is trained on s1K, s1K-1.1  (Muennighoff et al., 2025), and 3k samples drawn from the Mixture-of-Thoughts dataset of open-r1  (HuggingFace, 2025). Across the four reasoning benchmarks, our approach achieves relative improvements of 39%, 64%, and 83% over standard SFT.
Moreover, we observe that these gains persist through subsequent RL training stages. Compared with models cold-started using SFT, those initialized with WeFT achieve a relative performance improvement of 49% in later RL training. In addition, ablation studies further confirm that both the theoretically derived weighted SFT loss and the entropy-based weighting scheme are indispensable for achieving the observed performance improvements. Finally, the time efficiency analysis shows that WeFT introduces only minimal computational overhead, with merely a 24% increase in training time compared to SFT.
@@
  • We derive a theoretically grounded formulation of the weighted SFT loss from the perspective of diffusion processes, ensuring consistency with the underlying principles of diffusion-based generation.
  • We identify token-level predictive entropy as a reasonable and effective metric for capturing token importance, and we argue that tokens with higher entropy should be trained more frequently to enhance the model’s ability to handle planning and reasoning.
  • Building upon these insights, we propose WeFT, a weighted SFT algorithm that consistently outperforms the standard SFT approach across multiple training sets and benchmarks. Ablation studies further validate the necessity and effectiveness of each design choice in WeFT.
@@ 2 Preliminaries: Continuous-Time Discrete Diffusion Model
Following  (Ou et al., 2025), we model dLLMs as a Continuous-Time Discrete Diffusion Model (Sun et al., 2023). Intuitively, in each infinitesimal time interval $\Delta t$, every token has a certain probability of being masked.
@@
$p_{t+\Delta t|t}(y|x)=\left\{\begin{array}[]{ll}Q_{t}(x,y)\Delta t+o(t)&\text{if }x\neq y\\ 1+Q_{t}(x,x)\Delta t+o(t)&\text{if }x=y\end{array}\right$. (2)
@@ 3.1 Overview
We propose WeFT (Weighted Entropy-driven Fine-Tuning), a weighted SFT method designed to focus training on tokens that carry more information. In practice, we quantify token importance using the square root of token entropy.
As shown in Figure 1(a), the original SFT approach masks all tokens with the same probability $t$, and each token contributes equally to the loss function with a weight of $\frac{1}{t}$. In contrast, Figure 1(b) illustrates the WeFT pipeline. In each training step, we perform two forward passes. In the first pass, the entire answer is masked, and the masking rate $\beta_{i}$ for each token is estimated based on its predictive entropy. In the second pass, each token is independently masked with probability $t_{i}$ (computed from $\beta_{i}$), and its training weight is assigned as $\tfrac{1}{t_{i}}$. This design ensures that tokens with higher entropy are more likely to be masked and trained more frequently, thereby improving the model’s ability to capture complex dependencies.
The derivation of the loss function is presented in Section 3.2, while the details of the weighting scheme are provided in Section $3.$ A detailed description of the implementation of WeFT can be found in Appendix C.
3.2 Weighted SFT Loss Function Derivation
As noted earlier, prior work typically assumes that all tokens are masked at the same rate, leading to a $Q$ matrix that takes the form of an absorbing matrix composed of $1$ and 1-1. In contrast, we assign each token a distinct masking rate $\beta$, from which we derive a weighted SFT formulation. Importantly, the resulting SFT loss remains consistent with the fundamental properties of diffusion, since the $Q$ matrix is still well-defined under this construction. Namely, we now set $Q$ as:
@@ 3.2 Weighted SFT Loss Function Derivation
Based on the newly defined form of the $Q$ matrix, and following the derivation in  (Ou et al., 2025), we obtain the modified formulation of the SFT loss. The detailed derivation is provided in Appendix A.
{restatable}
theoremmaintheorem Assuming the $Q$ matrix takes the form given in Equation 9, let the initial sequence be $x_{0}$ and the sequence at time $t$ be $x_{t}$. Under this setting, the $i$-th token is masked with probability $t_{i}=1-(1-t)^{\tfrac{\beta_{x^{i}}}{\beta_{\text{ref}}}}$, where $\beta_{x^{i}}$ denotes the masking rate of the $i$-th token, and $\beta_{\text{ref}}$ is a specified reference masking rate. Moreover, the Weighted SFT loss can be derived as follows:
$L=\sum_{i}\mathbb{E}_{t_{i}}\left[\mathbf{1}\!\left[x_{t_{i}}^{i}=\mathbf{M}\right]\frac{1}{t_{i}}\log\big(x_{0}^{i}\mid x_{t}\big)\right]$ (10)
Intuitively, under the weighted SFT loss, a token $x_{i}$ with a larger masking rate $\beta_{x^{i}}$ corresponds to a larger $t_{i}$, making it more likely to be masked and subsequently learned. Here, $\beta_{\mathrm{ref}}$ can be any reference value. Empirically, we set $\beta_{\mathrm{ref}}$ to the mean of all $\beta$ values to enhance numerical stability.
@@ 3.3 Motivation of Using Entropy as the Masking Rate Metric
Having established the form of the weighted SFT loss, we next investigate how to assign masking rate to tokens. First, we require that the chosen metric does not introduce excessive computational overhead. In particular, it is infeasible to compute such a metric for every token individually, and we can only afford $O(1)$ additional forward passes in total. For this reason, we consider masking all the answers and then examining the distribution of the model’s output logits $z_{i}$ to be a practical choice. Specifically, given a problem $x$ and its answer $y$, we mask the entire answer and compute
@@
We then determine the masking rate $\beta_{i}$ of the $i$-th token. Based on our analysis, the entropy of the distribution of $z_{i}$ serves as an effective metric. As shown in Figure 2, we visualize the entropy distribution among the 100 most frequent tokens during training, contrasting tokens with higher versus lower entropy. It can be observed that tokens with higher entropy tend to carry richer information and play a more critical role in shaping the structure of generated responses. For example, words such as “first” and “second” exhibit higher entropy, often signaling logical or sequential structures, whereas words like “all” and “such” show lower entropy and function primarily as auxiliary words, contributing less to semantic content.
Accordingly, we define the masking rate $\beta_{i}$ as
$\beta_{i}=H(\mathrm{softmax}(z_{i})).$ (12)
However, we empirically find that the raw entropy values exhibit considerable variance, which may destabilize training. To mitigate this, we instead use the square root of entropy, which yields more stable optimization in practice. This design choice is further validated by experiments in Section  4.4.3.
@@
Frequent tokens with high entropy
Frequent tokens with low entropy
Figure 2: Visualization of high-frequency tokens with different entropy levels. Tokens with higher entropy tend to carry more information and play a crucial role in shaping the structure of generated responses.
@@ 4 Experiment
In this section, we first describe the experimental setup, followed by a comparison of the performance of WeFT and the standard SFT algorithm across multiple training datasets and benchmarks. We then analyze the performance of models trained with WeFT during subsequent RL training, conduct ablation studies to validate the necessity of each design component in WeFT, and finally provide a runtime comparison between WeFT and SFT under identical settings.
@@ 4.1 Experimental Setup
We adopt LLaDA-8B-Instruct  (Nie et al., 2025) as the base model and apply LoRA  (Hu et al., 2022) for fine-tuning. For the training datasets, we use s1K, s1K-1.1, and openr1-3k to evaluate the performance of WeFT and SFT. Among them, s1K and s1K-1.1 are reasoning datasets proposed by (Muennighoff et al., 2025), while openr1-3k is constructed by sampling 3k examples from the Mixture-of-Thoughts dataset of open-r1  (HuggingFace, 2025), with sequence lengths less than 2.5k under the open-r1 encoder. The evaluation is conducted on the following benchmarks:
@@
The detailed training and evaluation hyperparameters are provided in Appendix B.

### v2

# GIFT: Guided Importance-Aware Fine-Tuning for Diffusion Language Models

@@ Abstract
Diffusion models have recently shown strong potential in language modeling, offering faster generation compared to traditional autoregressive approaches. However, applying supervised fine-tuning (SFT) to diffusion models remains challenging, as they lack precise probability estimates at each denoising step. While the diffusion mechanism enables the model to reason over entire sequences, it also makes the generation process less predictable and often inconsistent. This highlights the importance of controlling key tokens that guide the direction of generation. To address this issue, we propose GIFT, an importance-aware finetuning method for diffusion language models, where tokens are assigned different importance weights based on their entropy. Derived from diffusion theory, GIFT delivers substantial gains: across diverse settings including different mainstream training datasets ranging from 1k to 10k in size, utilizing LoRA or full parameter fine-tuning, and training on base or instruct models, GIFT consistently achieves superior overall performance compared to standard SFT on four widely used reasoning benchmarks (Sudoku, Countdown, GSM8K, and MATH-500).
GIFT: Guided Importance-Aware Fine-Tuning for Diffusion Language Models
Guowei Xu†, Wenxin Xu†, Jiawang Zhao, Kaisheng Ma* Institute for Interdisciplinary Information Sciences Tsinghua University †Equal Contribution. Corresponding Author: kaisheng@tsinghua.edu.cn
@@ 1 Introduction
Recently, the landscape of large language modeling has been reshaped by the emergence of diffusion-based approaches  (Nie et al., 2025; Ou et al., 2025; Yang et al., 2025), which offer a fundamentally different perspective on sequence generation compared to autoregressive (AR) methods  Radford et al. (2018, 2019); Brown et al. (2020). Rather than relying on strictly sequential decoding, diffusion large language models (dLLMs) employ iterative refinement procedures that enable efficient parallel generation and open new possibilities for scaling language models.
@@
Specifically, when designing the SFT objective, Nie et al., 2025 model the generation process of dLLMs as a continuous-time diffusion process, where the diffusion timestep $t$ ranges within $[0,1]$. Here, $t=0$ denotes a completely original, non-masked sequence, and $t=1$ denotes a fully masked sequence. As illustrated in Figure 1(a), let the original response of the model be $x_{0}$. At a given timestep $t$, each token is masked with probability $t$, resulting in a partially masked response $x_{t}$. The training objective of dLLMs is then to enable the model to reconstruct the masked parts of $x_{0}$ given the partially masked input $x_{t}$. Based on this formulation, the loss is defined as follows:
$L=-\sum_{i}\mathbb{E}_{t}\left[\mathbf{1}[x_{t}^{i}=\mathbf{M}]\,\frac{1}{t}\log p\big(x_{0}^{i}\mid x_{t}\big)\right]$ (1)
@@
To address this discrepancy, we introduce GIFT, a guided importance-aware finetuning mechanism based on predictive uncertainty. Specifically, we find that the entropy of the model’s predictive distribution over tokens serves as a reliable metric for token importance at the current training stage. Tokens associated with higher entropy correspond to positions where the model exhibits greater uncertainty in generation. By prioritizing these high-entropy tokens during training via an adaptive importance weighting strategy applied to the SFT loss, we can concentrate computation and importance on tokens where the model is currently more uncertain, thereby improving the overall training effectiveness.
Figure 1: (a) The SFT pipeline. A timestep $t$ is uniformly sampled from $[0,1]$, and each token is masked independently with probability $t$. The training objective is to predict the masked tokens accurately based on the unmasked ones. (b) The GIFT pipeline. In each training step, we perform two forward passes. During the first forward pass, we mask the entire answer and estimate the masking rate $\beta_{i}$ for each token by computing its predictive entropy. In the second forward pass, the $i$-th token is masked with probability $t_{i}$ (computed from $\beta_{i}$), and its training weight is set to $\tfrac{1}{t_{i}}$. Tokens with higher entropy are more likely to be masked and thus receive stronger training signals.
Based on the fundamental principles of diffusion models, we derive an importance-weighted SFT loss formulation that is theoretically consistent with diffusion dynamics. Empirically, we evaluate this approach on four widely used benchmarks, Sudoku (Arel, 2025), Countdown (Pan et al., 2025), GSM8K (Cobbe et al., 2021), and MATH500 (Lightman et al., 2024), by comparing models fine-tuned with the standard SFT algorithm against those fine-tuned with GIFT under identical training data. Specifically, our method is trained using LoRA (Hu et al., 2022) on s1K, s1K-1.1 (Muennighoff et al., 2025), and 3k samples drawn from the Mixture-of-Thoughts dataset of open-r1 (HuggingFace, 2025), and using full parameter fine-tuning on 10k samples drawn from Tulu3 (Lambert et al., 2025). Across these diverse training settings, our approach consistently achieves superior overall performance compared to standard SFT.
In addition, ablation studies further confirm that both the theoretically derived importance-weighted SFT loss and the entropy-based importance-weighting scheme are indispensable for achieving the observed performance improvements. Finally, the time efficiency analysis shows that GIFT incurs only minimal overhead during LoRA fine-tuning, and imposes no computational overhead during full-parameter fine-tuning, even achieving faster speeds than standard SFT.
@@
  • We derive a theoretically grounded formulation of the importance-weighted SFT loss from the perspective of diffusion processes, ensuring consistency with the underlying principles of diffusion-based generation.
  • We identify token-level predictive entropy as a reasonable and effective metric for capturing token importance, and we demonstrate both theoretically and empirically that tokens with higher entropy should be emphasized during training to enhance the effectiveness and stability of the training process.
  • Building upon these insights, we propose GIFT, an importance-weighted finetuning algorithm that consistently achieves superior overall performance compared to the standard SFT approach across multiple training sets and benchmarks. Ablation studies further validate the necessity and effectiveness of each design choice in GIFT. Time efficiency analysis demonstrates that GIFT incurs minimal training time overhead compared to SFT.
@@ 2 Preliminaries: Continuous-Time Discrete Diffusion Model
Following  Ou et al., 2025, we model diffusion language models as a Continuous-Time Discrete Diffusion Model (Sun et al., 2023). Intuitively, in each infinitesimal time interval $\Delta t$, every token has a certain probability of being masked.
@@
$p_{t+\Delta t|t}(y|x)=\left\{\begin{aligned} &Q_{t}(x,y)\Delta t+o(\Delta t),\\ &\quad\text{if }x\neq y\\ &1+Q_{t}(x,x)\Delta t+o(\Delta t),\\ &\quad\text{if }x=y\end{aligned}\right$. (2)
@@
Frequent tokens with high entropy
Frequent tokens with low entropy
Figure 2: Visualization of high-frequency tokens with different entropy levels. Tokens with higher entropy correspond to those where the model exhibits greater uncertainty, and are learned more frequently during training.
@@ 3.1 Overview
We propose GIFT (Guided Importance-Aware Fine-Tuning), an importance-weighted finetuning method designed to assign higher importance weights to tokens where the current model exhibits greater uncertainty. As we will explain later, in practice, we quantify token importance using the square root of token entropy.
As shown in Figure 1(a), the original SFT approach masks all tokens with the same probability $t$, and each token contributes equally to the loss function with a weight of $\frac{1}{t}$. In contrast, Figure 1(b) illustrates the GIFT pipeline. In each training step, we perform two forward passes. In the first pass, the entire answer is masked, and the masking rate $\beta_{i}$ for each token is estimated based on its predictive entropy. In the second pass, each token is independently masked with probability $t_{i}$ (computed from $\beta_{i}$), and its training weight is assigned as $\tfrac{1}{t_{i}}$. This design ensures that tokens with higher entropy are more likely to be masked and trained more frequently, thereby improving the effectiveness and stability of the training process.
The derivation of the loss function is presented in Section 3.2, while the details and motivation of the importance weighting scheme are provided in Section 3.3 and Appendix D. A comprehensive description of the implementation of GIFT can be found in Appendix B.
3.2 Importance-aware Loss Derivation
As noted earlier, prior work typically assumes that all tokens are masked at the same rate, leading to a $Q$ matrix that takes the form of an absorbing matrix composed of $1$ and 1-1. In contrast, we assign each token a distinct masking rate $\beta$, from which we derive an importance-weighted finetuning formulation. Importantly, the resulting SFT loss remains consistent with the fundamental properties of diffusion, since the $Q$ matrix is still well-defined under this construction.
Namely, we now set $Q$ as:
@@ 3.2 Importance-aware Loss Derivation
Based on the newly defined form of the $Q$ matrix, and following the derivation in  Ou et al., 2025, we obtain the modified formulation of the SFT loss. The detailed derivation is provided in Appendix A.
Theorem 1.
Assuming the $Q$ matrix takes the form given in Equation 9, let the initial sequence be $x_{0}$ and the sequence at time $t$ be $x_{t}$. Under this setting, the $i$-th token is masked with probability $t_{i}=1-(1-t)^{\tfrac{\beta_{x^{i}}}{\beta_{\text{ref}}}}$, where $\beta_{x^{i}}$ denotes the masking rate of the $i$-th token, and $\beta_{\text{ref}}$ is a specified reference masking rate. Moreover, the importance-weighted supervised finetuning loss can be derived as follows:
$L=-\sum_{i}\mathbb{E}_{t_{i}}\left[\mathbf{1}\!\left[x_{t_{i}}^{i}=\mathbf{M}\right]\frac{1}{t_{i}}\log p\big(x_{0}^{i}\mid x_{t}\big)\right]$ (10)
Intuitively, under the importance-weighted supervised finetuning loss, a token $x_{i}$ with a larger masking rate $\beta_{x^{i}}$ corresponds to a larger $t_{i}$, making it more likely to be masked and subsequently learned. Here, $\beta_{\mathrm{ref}}$ can be any reference value. Empirically, we set $\beta_{\mathrm{ref}}$ to the mean of all $\beta$ values to enhance numerical stability.
@@ 3.3 Motivation of Using Entropy as the Masking Rate Metric
Having established the form of the importance-weighted supervised finetuning loss, we next investigate how to assign masking rate to tokens. First, we require that the chosen metric does not introduce excessive computational overhead. In particular, it is infeasible to compute such a metric for every token individually, and we can only afford $O(1)$ additional forward passes in total. For this reason, we consider masking all the answers and then examining the distribution of the model’s output logits $z_{i}$ to be a practical choice. Specifically, given a problem $x$ and its answer $y$, we mask the entire answer and compute
@@
We then determine the masking rate $\beta_{i}$ of the $i$-th token. Based on our analysis, the entropy of the distribution of $z_{i}$ serves as an effective metric. First, in Appendix D, drawing inspiration from theoretical insights, we demonstrate that using entropy as an importance weight may help reduce the variance of the loss estimation. Furthermore, as shown in Figure 2, we empirically visualize the entropy distribution among the 100 most frequent tokens during training, contrasting tokens with higher versus lower entropy. It can be observed that tokens with higher entropy tend to carry richer information and play a more critical role in the model’s decision-making. For example, words such as “first” and “second” exhibit higher entropy, often signaling logical or sequential structures, whereas words like “all” and “such” show lower entropy and function primarily as auxiliary words, contributing less to semantic content.
Since we empirically find that using the raw entropy values may destabilize training (Section  4.3.3), we use the square root of entropy, which yields more stable optimization in practice:
@@ 4 Experiment
In this section, we first describe the experimental setup, followed by a comparison of the performance of GIFT and the standard SFT algorithm across multiple training datasets and benchmarks. We then conduct ablation studies to validate the necessity of each design component in GIFT, and finally provide a runtime comparison between GIFT and SFT under identical settings.
Model Sudoku Countdown GSM8K MATH500 Average
0-shot 0-shot 0-shot 0-shot
Base Instruct Model
LLaDA-8B-Instruct 5.5 16.0 76.7 32.4 32.6
s1K Dataset, LoRA Tuning
Instruct Model + SFT 4.6 (-0.9) 23.8 (+7.8) 78.8 (+2.1) 32.6 (+0.2) 34.9 (+2.3)
Instruct Model + GIFT 7.8 (+2.3) 24.6 (+8.6) 78.0 (+1.3) 33.0 (+0.6) 35.8 (+3.2)
s1K-1.1 Dataset, LoRA Tuning
Instruct Model + SFT 4.2 (-1.3) 21.5 (+5.5) 78.4 (+1.7) 30.8 (-1.6) 33.7 (+1.1)
Instruct Model + GIFT 6.8 (+1.3) 20.3 (+4.3) 78.1 (+1.4) 32.6 (+0.2) 34.4 (+1.8)
openr1-3k Dataset, LoRA Tuning
Instruct Model + SFT 6.3 (+0.8) 17.2 (+1.2) 77.3 (+0.6) 32.0 (-0.4) 33.2 (+0.6)
Instruct Model + GIFT 6.9 (+1.4) 17.6 (+1.6) 77.5 (+0.8) 32.8 (+0.4) 33.7 (+1.1)
Tulu3-10k Dataset, Full Paramter Tuning
Base Model + SFT 5.0 18.7 66.1 23.2 28.2
Base Model + GIFT 4.8 20.3 65.4 25.0 28.9
Table 1: Comparison of the performance of GIFT and SFT across different training datasets and benchmarks. The table shows that GIFT achieves substantial percentage improvements over SFT.
@@ 4.1 Experimental Setup
We adopt LLaDA-8B-Base and LLaDA-8B-Instruct (Nie et al., 2025) as base models and apply either LoRA fine-tuning (Hu et al., 2022) or full parameter fine-tuning. For the training datasets, we use s1K, s1K-1.1, openr1-3k, and Tulu3-10k to evaluate the performance of GIFT and SFT. Among them, s1K and s1K-1.1 are reasoning datasets proposed by Muennighoff et al., 2025, while openr1-3k is constructed by sampling 3k examples from the Mixture-of-Thoughts dataset of open-r1 (HuggingFace, 2025), with sequence lengths less than 2.5k under the open-r1 encoder. Tulu3-10k consists of 10k samples drawn from Tulu3 (Lambert et al., 2025), following the setting of Zhou et al., 2025. To verify that our algorithm is framework-agnostic, we employ the wd1  (Tang et al., 2025) framework for LoRA training, and the dllm (Zhou et al., 2025) framework for full-parameter fine-tuning.
The evaluation is conducted on the following benchmarks:
@@
The detailed training and evaluation hyperparameters and settings are provided in Appendix C.